OpenStudy (anonymous):

the arm of a steel crane is 52feet long.when it is fully extended the crane arm forms a 30 degree angle with a line parallel to the ground . how many feet above the ground is the top of the crane ?

OpenStudy (dumbcow):

let h be the height of the crane sin30 = opp/hyp sin30 = h/52 h = 52*sin30 --> sin30 = 1/2 (use unit circle or calc) h = 52*1/2 h = 26